2. In the simple regression model, show that Bo is consistent under the as- sumption SLR 4.PLZ HELP!
Answer #1

Assumption SLR.4 (Zero Conditional Mean )

E(u|x) = 0

Cov(x,e) = E(Xu) - E(X)E(u) = E(Xu) =E(E(XU|x))= E (xE(U|x)) = 0

Well look at β0-y--81z first. The law of large numbers says that y converges to E(y) Ao + gE(z) and if βι-+ A then βι x conv